Thursday, April 23: OWL Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans

  • General considerations
    • Kijun Park is the current Run Coordinator. Please contact directly via RC phone 757-876-1787
    • Follow the generic Instructions for shift takers
    • Shift leaders: Please make sure to fill out the beam time accounting form and write a "Shift Summary" log entry at the end of your shift regardless if there is beam or not.
    • In general, Hall technical staff should not be contacted by the shift crew - please call the RC first.
    • The target operator is required to supervise the cryo-target.
    • Make sure to use the DVCS CODA configuration for all runs (even when taking single arm HRS data).

Optics Calibration Data

    • Note: Keep an eye on the white board in CC more detail
    • 1. After finishing bulls-eye scan runs move to "optics" target
    • 2. For optics we will use only Right-HRS DAQ
    • 3. Make sure the BPM A and BPM D readings are at the nominal position (check white board)
    • 4. Set the prescale factors on Right HRS to ps1=1 and rest to zero
    • 5. Take one CODA run each for the following settings on Q1 current on Right HRS.
      • - We need at least 10 K good electron events (determined by an expert in the backroom)
        • Q1 current = 0 A
        • Q1 Current = 167.7375 A
    • 6. Change Right-HRS Q1 current to 153.36 A
    • 7. Take a CODA run with Right HRS. We need at least 100 K electron events (determined by an expert)